# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
$e(\Gamma)=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}\left[\left(y_{\ell+1}-\langle\Gamma,H_{Z}\rangle\right)^{2}\right],$ (SI.2.1)
where $(Z,y_{\ell+1})\sim\mathcal{P}_{\mathrm{test}}$ is a new sample drawn from the distribution of the test data set. Recall that $Z$ is the input embedding matrix defined in (4) in the main text, and $y_{\ell+1}$ denotes the missing value to be predicted. The goal of this section is to derive an expression for the generalization error $e(\Gamma)$.
@@
$\mathbb{E}_{w}\left[y_{\ell+1}\right]=0,\quad\text{and}\quad\mathbb{E}_{w}\left[H_{Z}\right]=0.$ (SI.2.2)
@@
$\mathbb{E}_{w}\left[y_{\ell+1}H_{Z}\right]=\frac{1}{d}w\begin{bmatrix}w^{\top},&1+\rho\end{bmatrix}$ (SI.2.3)
@@
$\mathbb{E}_{w}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]=\frac{(1+\rho)}{d}I_{d}\otimes\begin{bmatrix}\frac{d}{\ell}I_{d}+(1+\ell^{-1})(1+\rho)^{-1}ww^{\top}&(1+2\ell^{-1})w\\ (1+2\ell^{-1})w^{\top}&(1+2\ell^{-1})(1+\rho)\end{bmatrix}.$ (SI.2.4)
@@
Using the equivalent representations in (SI.7.1) and (SI.7.2), it is straightforward to verify the estimates of the first (conditional) moments in (SI.2.2). To show (SI.2.3), we note that
$H_{Z}=(d/\ell)z_{a}z_{b}^{\top},$ (SI.2.5)
@@
$z_{a}=M_{w}\begin{bmatrix}s\\ u\end{bmatrix}\qquad\text{and}\qquad z_{b}=\begin{bmatrix}M_{w}h\\ (\theta_{w}a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\end{bmatrix}.$ (SI.2.6)
Using the representation in (SI.7.2), we have
$\mathbb{E}_{w}\left[y_{\ell+1}H_{Z}\right]=(d/\ell)\mathbb{E}_{w}\left[y_{\ell+1}z_{a}\right]\mathbb{E}_{w}\left[z_{b}^{\top}\right].$ (SI.2.7)
Computing the expectations $\mathbb{E}_{w}\left[y_{\ell+1}z_{a}\right]$ and $\mathbb{E}_{w}\left[z_{b}^{\top}\right]$ then gives us (SI.2.3). Next, we show (SI.2.4). Since $z_{a}$ and $z_{b}$ are independent,
$\mathbb{E}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]=(d/\ell)^{2}\,\mathbb{E}\left[z_{a}z_{a}^{\top}\right]\otimes\mathbb{E}\left[z_{b}z_{b}^{\top}\right].$ (SI.2.8)
@@
$\mathbb{E}_{w}\left[z_{a}z_{a}^{\top}\right]=I_{d}$ (SI.2.9)
@@
$\mathbb{E}_{w}\left[M_{w}hh^{\top}M_{w}\right]=\frac{\ell}{d^{2}}\left[(1+\rho)I_{d}+\frac{(\ell+1)}{d}ww^{\top}\right].$ (SI.2.10)
@@
$\mathbb{E}_{w}\left[M_{w}h\left((a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\right)\right]=\frac{\ell(\ell+2)(1+\rho)}{d^{3}}w$ (SI.2.11)
@@
$\mathbb{E}_{w}\left[\left((a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\right)^{2}\right]=\frac{\ell(\ell+2)(1+\rho)^{2}}{d^{3}}.$ (SI.2.12)
Combining (SI.2.10), (SI.2.11), and (SI.2.12), we have
$\mathbb{E}\left[z_{b}z_{b}^{\top}\right]=\frac{(\ell/d)^{2}(1+\rho)}{d}\begin{bmatrix}\frac{d}{\ell}I_{d}+(1+\ell^{-1})(1+\rho)^{-1}ww^{\top}&(1+2\ell^{-1})w\\ (1+2\ell^{-1})w^{\top}&(1+2\ell^{-1})(1+\rho)\end{bmatrix}.$ (SI.2.13)
Substituting (SI.2.9) and (SI.2.13) into (SI.2.4), we reach the formula in (SI.2.4). ∎
@@
$e(\Gamma)$ $=\frac{1+\rho}{d}\operatorname{tr}\left(\Gamma\begin{bmatrix}\frac{d}{\ell}I_{d}+(1+\ell^{-1})(1+\rho)^{-1}R_{\mathrm{test}}&(1+2\ell^{-1})b_{\mathrm{test}}\\ (1+2\ell^{-1})b_{\mathrm{test}}^{\top}&(1+2\ell^{-1})(1+\rho)\end{bmatrix}\Gamma^{\top}\right)$ (SI.2.14)
@@
$b_{\mathrm{test}}\coloneqq\mathbb{E}_{w\sim\mathcal{P}_{\mathrm{test}}}\left[w\right]\quad\text{and}\quad R_{\mathrm{test}}\coloneqq\mathbb{E}_{w\sim\mathcal{P}_{\mathrm{test}}}\left[ww^{\top}\right].$ (SI.2.15)
@@
We use $w\sim\mathcal{P}_{\mathrm{test}}$ to indicate that $w$ is sampled from the task vector distribution associated with $\mathcal{P}_{\mathrm{test}}$. Recall our discussions in Section II.6. For the ICL task, $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. It is straightforward to check that, in this case,
[…]
@@
[…]
@@
Recall the definition of the generalization error in (SI.2.1). We start by writing
$e(\Gamma)=\operatorname{vec}(\Gamma)^{\top}\mathbb{E}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]\operatorname{vec}(\Gamma)-2\operatorname{vec}(\Gamma)^{\top}\operatorname{vec}(\mathbb{E}\left[y_{N+1}H_{Z}\right])+\mathbb{E}\left[y_{\ell+1}^{2}\right],$ (SI.2.18)
where $H_{Z}$ is a matrix in the form of (17) and $H_{Z}$ is independent of $\Gamma$. Since $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\rho)$ denoting the noise, it is straightforward to check that
$\mathbb{E}\left[y_{\ell+1}^{2}\right]=1+\rho.$ (SI.2.19)
Using the moment estimate (SI.2.4) in Lemma 1 and the identity (SI.1.1), we have
$\operatorname{vec}(\Gamma)^{\top}\mathbb{E}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]\operatorname{vec}(\Gamma)$ (SI.2.20)
@@
Moreover, by (SI.2.3),
$\operatorname{vec}(\Gamma)^{\top}\operatorname{vec}\left(\mathbb{E}\left[y_{\ell+1}H_{Z}\right]\right)=\frac{1}{d}\operatorname{tr}\left(\Gamma\begin{bmatrix}R_{\mathrm{test}}\\ (1+\rho)b_{\mathrm{test}}^{\top}\end{bmatrix}\right).$ (SI.2.21)
@@
$e(\Gamma)=\frac{1}{d}\operatorname{tr}\left(\Gamma B_{\mathrm{test}}\Gamma^{\top}\right)-\frac{2}{d}\operatorname{tr}\left(\Gamma A_{\mathrm{test}}^{\top}\right)+(1+\rho)+\mathcal{E},$ (SI.2.22)
@@
$A_{\mathrm{test}}\coloneqq\begin{bmatrix}R_{\mathrm{test}}&(1+\rho)b_{\mathrm{test}}\end{bmatrix},$ (SI.2.23)
$B_{\mathrm{test}}\coloneqq\begin{bmatrix}\frac{1}{\alpha}(1+\rho)I_{d}+R_{\mathrm{test}}&(1+\rho)b_{\mathrm{test}}\\ (1+\rho)b_{\mathrm{test}}^{\top}&(1+\rho)^{2}\end{bmatrix},$ (SI.2.24)
and $R_{\mathrm{test}},b_{\mathrm{test}}$ are as defined in (SI.2.15). Moreover, $\mathcal{E}$ denotes an “error” term such that
$\lvert\mathcal{E}\rvert\leq\frac{C_{\alpha,\rho}\max\left\{\lVert R_{\mathrm{test}}\rVert_{\mathsf{op}},\lVert b_{\mathrm{test}}\rVert,1\right\}\left(\lVert\Gamma\rVert_{\mathsf{F}}^{2}/d\right)}{d},$ (SI.2.25)
@@
$\Delta=\begin{bmatrix}\frac{d}{\ell}(1+\rho)I_{d}+(1+\ell^{-1})R_{\mathrm{test}}&(1+2\ell^{-1})(1+\rho)b_{\mathrm{test}}\\ (1+2\ell^{-1})(1+\rho)b_{\mathrm{test}}^{\top}&(1+2\ell^{-1})(1+\rho)^{2}\end{bmatrix}-B_{\mathrm{test}}.$ (SI.2.26)
@@
$\mathcal{E}$ $=\frac{1}{d}\operatorname{tr}\left(\Gamma\Delta\Gamma^{\top}\right)$ (SI.2.27)
$=\frac{1}{d}\operatorname{vec}(\Gamma)^{\top}(I_{d}\otimes\Delta)\operatorname{vec}(\Gamma)$ (SI.2.28)
$\leq\lVert\Delta\rVert_{\mathsf{op}}\frac{\lVert\Gamma\rVert_{\mathsf{F}}^{2}}{d}.$ (SI.2.29)
The bound in (SI.2.25) follows from the estimate that $\lVert\Delta\rVert_{\mathsf{op}}\leq C_{\alpha,\rho}\max\left\{\lVert R_{\mathrm{test}}\rVert_{\mathsf{op}},\lVert b_{\mathrm{test}}\rVert,1\right\}/d$. ∎
@@
Consider the optimal weight matrix $\Gamma^{\ast}$ obtained by solving the ridge regression problem in (20). Since $\Gamma^{\ast}$ is the optimal solution of (20), we must have
$\frac{n}{d}\lambda\lVert\Gamma^{\ast}\rVert_{\mathsf{F}}^{2}\leq\sum_{\mu\in[n]}(y_{\ell+1}^{\mu})^{2},$ (SI.2.30)
@@
$\frac{\lVert\Gamma^{\ast}\rVert_{\mathsf{F}}^{2}}{d}\leq\frac{\sum_{\mu\in[n]}(y_{\ell+1}^{\mu})^{2}}{\lambda\,n}.$ (SI.2.31)
By the law of large numbers, $\frac{\sum_{\mu\in[n]}y_{\mu}^{2}}{n}\to 1+\rho$ as $n\to\infty$. Thus, $\lVert\Gamma^{\ast}\rVert_{\mathsf{F}}^{2}/{d}$ is asymptotically bounded by the constant $(1+\rho)/\lambda$. Furthermore, it is easy to check that $\lVert R_{\mathrm{test}}\rVert_{\mathsf{op}}=\mathcal{O}(1)$ and $\lVert b_{\mathrm{test}}\rVert=\mathcal{O}(1)$ for both ICL [see (SI.2.16)] and IDG [see (SI.2.17)]. It then follows from Corollary 1 that the generalization error associated with the optimal parameers $\Gamma^{\ast}$ is asymptotically determined by the first three terms on the right-hand side of (SI.2.22).
SI.3 Analysis of Ridge Regression: Extended Resolvent Matrices

### v3

@@ Appendix B Moment Calculations and Generalization Errors
$e(\Gamma)=\mathbb{E}_{\mathcal{P}_{\mathrm{test}}}\left[\left(y_{\ell+1}-\langle\Gamma,H_{Z}\rangle\right)^{2}\right],$ (44)
where $(Z,y_{\ell+1})\sim\mathcal{P}_{\mathrm{test}}$ is a new sample drawn from the distribution of the test data set. Recall that $Z$ is the input embedding matrix defined by
and $y_{\ell+1}$ denotes the missing value to be predicted. The goal of this section is to derive an expression for the generalization error $e(\Gamma)$.
@@
$\mathbb{E}_{w}\left[y_{\ell+1}\right]=0,\quad\text{and}\quad\mathbb{E}_{w}\left[H_{Z}\right]=0.$ (48)
@@
$\mathbb{E}_{w}\left[y_{\ell+1}H_{Z}\right]=\frac{1}{d}w\begin{bmatrix}w^{\top},&1+\rho\end{bmatrix}$ (49)
@@
$\mathbb{E}_{w}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]=\frac{(1+\rho)}{d}I_{d}\otimes\begin{bmatrix}\frac{d}{\ell}I_{d}+(1+\ell^{-1})(1+\rho)^{-1}ww^{\top}&(1+2\ell^{-1})w\\ (1+2\ell^{-1})w^{\top}&(1+2\ell^{-1})(1+\rho)\end{bmatrix}$. (50)
@@
Using the equivalent representations in (233) and (234), it is straightforward to verify the estimates of the first (conditional) moments in (48). To show (49), we note that
$H_{Z}=(d/\ell)z_{a}z_{b}^{\top},$ (51)
@@
$z_{a}=M_{w}\begin{bmatrix}s\\ u\end{bmatrix}\qquad\text{and}\qquad z_{b}=\begin{bmatrix}M_{w}h\\ (\theta_{w}a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\end{bmatrix}$. (52)
Using the representation in (234), we have
$\mathbb{E}_{w}\left[y_{\ell+1}H_{Z}\right]=(d/\ell)\mathbb{E}_{w}\left[y_{\ell+1}z_{a}\right]\mathbb{E}_{w}\left[z_{b}^{\top}\right].$ (53)
Computing the expectations $\mathbb{E}_{w}\left[y_{\ell+1}z_{a}\right]$ and $\mathbb{E}_{w}\left[z_{b}^{\top}\right]$ then gives us (49). Next, we show (50). Since $z_{a}$ and $z_{b}$ are independent,
$\mathbb{E}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]=(d/\ell)^{2}\,\mathbb{E}\left[z_{a}z_{a}^{\top}\right]\otimes\mathbb{E}\left[z_{b}z_{b}^{\top}\right].$ (54)
@@
$\mathbb{E}_{w}\left[z_{a}z_{a}^{\top}\right]=I_{d}$ (55)
@@
$\mathbb{E}_{w}\left[M_{w}hh^{\top}M_{w}\right]=\frac{\ell}{d^{2}}\left[(1+\rho)I_{d}+\frac{(\ell+1)}{d}ww^{\top}\right].$ (56)
@@
$\mathbb{E}_{w}\left[M_{w}h\left((a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\right)\right]=\frac{\ell(\ell+2)(1+\rho)}{d^{3}}w$ (57)
@@
$\mathbb{E}_{w}\left[\left((a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\right)^{2}\right]=\frac{\ell(\ell+2)(1+\rho)^{2}}{d^{3}}.$ (58)
Combining (56), (57), and (58), we have
$\mathbb{E}\left[z_{b}z_{b}^{\top}\right]=\frac{(\ell/d)^{2}(1+\rho)}{d}\begin{bmatrix}\frac{d}{\ell}I_{d}+(1+\ell^{-1})(1+\rho)^{-1}ww^{\top}&(1+2\ell^{-1})w\\ (1+2\ell^{-1})w^{\top}&(1+2\ell^{-1})(1+\rho)\end{bmatrix}$. (59)
Substituting (55) and (59) into (50), we reach the formula in (50). ∎
@@
$e(\Gamma)$ $=\frac{1+\rho}{d}\operatorname{tr}\left(\Gamma\begin{bmatrix}\frac{d}{\ell}I_{d}+(1+\ell^{-1})(1+\rho)^{-1}R_{\mathrm{test}}&(1+2\ell^{-1})b_{\mathrm{test}}\\ (1+2\ell^{-1})b_{\mathrm{test}}^{\top}&(1+2\ell^{-1})(1+\rho)\end{bmatrix}\Gamma^{\top}\right)$ (60)
@@
$b_{\mathrm{test}}\coloneqq\mathbb{E}_{w\sim\mathcal{P}_{\mathrm{test}}}\left[w\right]\quad\text{and}\quad R_{\mathrm{test}}\coloneqq\mathbb{E}_{w\sim\mathcal{P}_{\mathrm{test}}}\left[ww^{\top}\right].$ (61)
@@
We use $w\sim\mathcal{P}_{\mathrm{test}}$ to indicate that $w$ is sampled from the task vector distribution associated with $\mathcal{P}_{\mathrm{test}}$. For the ICL task, $w\sim\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. It is straightforward to check that, in this case,
$\textrm{(ICL)}\mathrel{\mathop{\ordinarycolon}}\qquad b_{\mathrm{test}}=0\quad\text{and}\quad R_{\mathrm{test}}=I_{d}.$ (62)
@@
$\text{(IDG)}\mathrel{\mathop{\ordinarycolon}}\qquad b_{\mathrm{test}}=\frac{1}{k}\sum_{i}w_{i}\quad\text{and}\quad R_{\mathrm{test}}=\frac{1}{k}\sum_{i\in[k]}w_{i}w_{i}^{\top},$ (63)
@@
Recall the definition of the generalization error in (44). We start by writing
$e(\Gamma)=\operatorname{vec}(\Gamma)^{\top}\mathbb{E}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]\operatorname{vec}(\Gamma)-2\operatorname{vec}(\Gamma)^{\top}\operatorname{vec}(\mathbb{E}\left[y_{N+1}H_{Z}\right])+\mathbb{E}\left[y_{\ell+1}^{2}\right],$ (64)
where $H_{Z}$ is a matrix in the form of
and $H_{Z}$ is independent of $\Gamma$. Since $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\rho)$ denoting the noise, it is straightforward to check that
$\mathbb{E}\left[y_{\ell+1}^{2}\right]=1+\rho.$ (66)
Using the moment estimate (50) in Lemma 1 and the identity (39), we have
$\operatorname{vec}(\Gamma)^{\top}\mathbb{E}\left[\operatorname{vec}(H_{Z})\operatorname{vec}(H_{Z})^{\top}\right]\operatorname{vec}(\Gamma)$ (67)
@@
Moreover, by (49),
[…]
@@
$e(\Gamma)=\frac{1}{d}\operatorname{tr}\left(\Gamma B_{\mathrm{test}}\Gamma^{\top}\right)-\frac{2}{d}\operatorname{tr}\left(\Gamma A_{\mathrm{test}}^{\top}\right)+(1+\rho)+\mathcal{E},$ (69)
@@
$A_{\mathrm{test}}\coloneqq\begin{bmatrix}R_{\mathrm{test}}&(1+\rho)b_{\mathrm{test}}\end{bmatrix}$, (70)
$B_{\mathrm{test}}\coloneqq\begin{bmatrix}\frac{1}{\alpha}(1+\rho)I_{d}+R_{\mathrm{test}}&(1+\rho)b_{\mathrm{test}}\\ (1+\rho)b_{\mathrm{test}}^{\top}&(1+\rho)^{2}\end{bmatrix}$, (71)
and $R_{\mathrm{test}},b_{\mathrm{test}}$ are as defined in (61). Moreover, $\mathcal{E}$ denotes an “error” term such that
$\mathinner{\!\left\lvert\mathcal{E}\right\rvert}\leq\frac{C_{\alpha,\rho}\max\left\{\mathinner{\!\left\lVert R_{\mathrm{test}}\right\rVert}_{\mathsf{op}},\mathinner{\!\left\lVert b_{\mathrm{test}}\right\rVert},1\right\}\left(\mathinner{\!\left\lVert\Gamma\right\rVert}_{\mathsf{F}}^{2}/d\right)}{d},$ (72)
@@
$\Delta=\begin{bmatrix}\frac{d}{\ell}(1+\rho)I_{d}+(1+\ell^{-1})R_{\mathrm{test}}&(1+2\ell^{-1})(1+\rho)b_{\mathrm{test}}\\ (1+2\ell^{-1})(1+\rho)b_{\mathrm{test}}^{\top}&(1+2\ell^{-1})(1+\rho)^{2}\end{bmatrix}-B_{\mathrm{test}}.$ (73)
@@
$\mathcal{E}$ $=\frac{1}{d}\operatorname{tr}\left(\Gamma\Delta\Gamma^{\top}\right)$ (74)
$=\frac{1}{d}\operatorname{vec}(\Gamma)^{\top}(I_{d}\otimes\Delta)\operatorname{vec}(\Gamma)$ (75)
$\leq\mathinner{\!\left\lVert\Delta\right\rVert}_{\mathsf{op}}\frac{\mathinner{\!\left\lVert\Gamma\right\rVert}_{\mathsf{F}}^{2}}{d}.$ (76)
The bound in (72) follows from the estimate that $\mathinner{\!\left\lVert\Delta\right\rVert}_{\mathsf{op}}\leq C_{\alpha,\rho}\max\left\{\mathinner{\!\left\lVert R_{\mathrm{test}}\right\rVert}_{\mathsf{op}},\mathinner{\!\left\lVert b_{\mathrm{test}}\right\rVert},1\right\}/d$. ∎
@@
Consider the optimal weight matrix $\Gamma^{\ast}$ obtained by solving the ridge regression problem in
$\Gamma^{\ast}$ $=\underset{\Gamma}{\arg\,\min}\,\sum_{\mu=1}^{n}\left(y_{\ell+1}^{\mu}-\langle\Gamma,H_{Z^{\mu}}\rangle\right)^{2}+\frac{n}{d}\lambda\mathinner{\!\left\lVert\Gamma\right\rVert}_{\mathrm{F}}^{2}\,.$ (77)
Since $\Gamma^{\ast}$ is the optimal solution of (77), we must have
$\frac{n}{d}\lambda\mathinner{\!\left\lVert\Gamma^{\ast}\right\rVert}_{\mathsf{F}}^{2}\leq\sum_{\mu\in[n]}(y_{\ell+1}^{\mu})^{2},$ (78)
@@
$\frac{\mathinner{\!\left\lVert\Gamma^{\ast}\right\rVert}_{\mathsf{F}}^{2}}{d}\leq\frac{\sum_{\mu\in[n]}(y_{\ell+1}^{\mu})^{2}}{\lambda\,n}.$ (79)
By the law of large numbers, $\frac{\sum_{\mu\in[n]}y_{\mu}^{2}}{n}\to 1+\rho$ as $n\to\infty$. Thus, $\mathinner{\!\left\lVert\Gamma^{\ast}\right\rVert}_{\mathsf{F}}^{2}/{d}$ is asymptotically bounded by the constant $(1+\rho)/\lambda$. Furthermore, it is easy to check that $\mathinner{\!\left\lVert R_{\mathrm{test}}\right\rVert}_{\mathsf{op}}=\mathcal{O}(1)$ and $\mathinner{\!\left\lVert b_{\mathrm{test}}\right\rVert}=\mathcal{O}(1)$ for both ICL [see (62)] and IDG [see (63)]. It then follows from Corollary 1 that the generalization error associated with the optimal parameers $\Gamma^{\ast}$ is asymptotically determined by the first three terms on the right-hand side of (69).
Appendix C Analysis of Ridge Regression: Extended Resolvent Matrices